Find the equation, in standard form, of the line perpendicular to -5x+3y=3 that passes through the points (-8,-9). Write the equations in standard form.

Answer by TimothyLamb(4379)

---
-5x + 3y = 3
3y = 5x + 3
y = (5/3)x + 1
---
perp lines have negative reciprocal slopes, so the perp slope = -3/5 = -0.6
---
perp line contains (-8,-9)
---
to solve the linear equation, copy and paste this (given point and perp slope):
-8,-9,-0.6
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (-23, 0)
y-intercept = (0, -13.8)
slope = -0.6
---
answer:
the slope-intercept form:
y = -0.6x - 13.8
the standard form:
0.6x + y = -13.8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php